# Benford behavior resulting from stick and box fragmentation processes

### v2

@@ 3.8. Evaluation of main term $\mathcal{M}_{\textup{main},N}(a,b)$
$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (A.1)
@@
$\ =\ \frac{k(N)!\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!-k(N)!\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}$
@@
$\cdot\frac{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!-\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}$
$\ =\ \binom{k(N)}{k_{1,\ell}}\left(1-\frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}\right).$ (A.2)
@@
$\alpha_{\ell,N}\ :=\ \frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}.$ (A.3)
@@
$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}$ $\ \leq\ \alpha_{\ell,N}\ \leq\left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (A.4)
@@
$0\ \leq\ \frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\ \leq\ \frac{4\ell N^{\delta}}{k(N)}\ \leq\ \frac{2N^{\delta}}{\sqrt{k(N)}}.$ (A.5)
@@
$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/10}}{N^{\epsilon/2}}\ =\ 2N^{-2\epsilon/5}\ =\ O(N^{-2\epsilon/5}).$ (A.6)
@@
$\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\ =\ O(N^{-2\epsilon/5}).$ (A.7)
@@
$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}\ \leq\ \alpha_{\ell,N}\ \leq\ \left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (A.9)
@@
$\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}$ $\ \leq\ \alpha_{\ell,N}-1\ \leq\ \sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{j}.$ (A.10)
@@
$\ =\ \frac{2N^{2\delta}}{\sqrt{k(N)}}\cdot\frac{1-\left(\frac{2N^{2\delta}}{\sqrt{k(N)}}\right)^{N^{\delta}}}{1-\frac{2N^{2\delta}}{\sqrt{k(N)}}},$ (A.11)
@@
$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/5}}{N^{\epsilon/2}}\ =\ 2N^{-3\epsilon/10}.$ (A.12)
@@
$\ =\ O\left(N^{-3\epsilon/10}\right).$ (A.13)
@@
$\left|\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (A.14)
Applying (A) and (A.14) to (A), we get
$\left|\alpha_{\ell,N}-1\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (A.15)
Thus, substituting the estimate (A.15) back to (A) gives us
$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-3\epsilon/10}\right).$ (A.16)
@@ Appendix B Case for $k_{S}<d$
$D_{1}(y)$ $\ :=\ \sum_{\begin{subarray}{c}S\subset[d]\\ S\neq[d]\end{subarray}}\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
$\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (B.1)
@@
$\quad\quad\quad\cdot\left(\prod_{i\in[k_{S}+1;d-1]}\varphi(z_{m+i-d})\right)\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right).$ (B.2)
@@
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
@@
$\quad\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
$\quad\quad\quad\cdot A(z_{m+k_{S}-d})\cdot\left(\prod_{i=k_{S}+1}^{d-1}\varphi(z_{m+i-d})\right)\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (B.3)
To bound (B), we first give some estimate on the following
@@
$\quad\ \ll\ \frac{1}{\sqrt{2\pi(d-k_{S})}}e^{-\frac{\left(y-\sum_{i=1}^{k_{S}}z_{m+i-d}\right)^{2}}{2(d-k_{S})}},$ (B.4)
@@
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{k_{S}-1}\int_{z_{m}+(j-1)-d}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
@@
$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\underbrace{\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\cdots\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}}_{k_{S}-1}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
@@
$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\prod_{i\in S\setminus[k_{S}+1;d]}\left(\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\varphi(z_{m+i-d})dz_{m+i-d}\right)\cdot\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}\left(\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}A(z_{m+i-d})dz_{m+i-d}\right)$
$\quad\ \ll\ N^{-1/2-\left(1+\left|[d]\setminus(S\cup\{k_{S}\})\right|\right)\delta}\ \ll\ N^{-1/2-\delta},$ (B.5)
@@
Since the number of proper subsets $S$ of $[d]$ is finite, then when $k_{S}<d$, we have that $D_{1}(y)\ll N^{-1/2-\delta}$ by definition of $D_{1}(y)$ in (B).
@@ Appendix C Case for $d>2$
In this section, we want to obtain the following estimate on $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ when $d>2$
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
@@
$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (C.1)
@@
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
$\quad\ \ll\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg{)}dz_{m-2}\cdots dz_{m+1-d}.$ (C.2)
Since $y$ is an upper bound on $\sum_{i=1}^{d}z_{m+i-d}$ and $z_{m-2}\leq z_{m-1},z_{m}$, then we have
$z_{m-2}\ \leq\ \frac{z_{m-1}+z_{m}}{2}\ \leq\ \frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}.$ (C.3)
When $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi((y-\sum_{i=1}^{d-2}z_{m+i-d})/2)$. Hence,
@@
$\quad\ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1},$ (C.4)
where in the last line we use the change of variable $z_{m-1}\mapsto-z_{m-1}+y-\sum_{i=1}^{d-2}z_{m+i-d}$. The integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard normal random variable, which is finite. Hence, when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$,
@@
$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right).$ (C.5)
@@
$y-\sum_{i=1}^{d-1}z_{m+i-d}\ \geq\ y-\sum_{i=1}^{d-2}z_{m+i-d}-\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\ =\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\ \geq\ 0.$ (C.6)
@@
$\quad\ =\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)-\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (C.7)
@@
$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)+\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (C.8)
@@
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
$\quad\ \ll\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
@@
$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d},$ (C.9)

### v3

@@ Appendix A Proof of Proposition 3
$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (152)
@@
$\cdot\frac{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!-\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}$
$\ =\ \binom{k(N)}{k_{1,\ell}}\left(1-\frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}\right).$ (153)
@@
$\alpha_{\ell,N}\ :=\ \frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}.$ (154)
@@
$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}$ $\ \leq\ \alpha_{\ell,N}\ \leq\left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (155)
@@
$0\ \leq\ \frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\ \leq\ \frac{4\ell N^{\delta}}{k(N)}\ \leq\ \frac{2N^{\delta}}{\sqrt{k(N)}}.$ (156)
@@
$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/10}}{N^{\epsilon/2}}\ =\ 2N^{-2\epsilon/5}\ =\ O(N^{-2\epsilon/5}).$ (157)
@@
$\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\ =\ O(N^{-2\epsilon/5}).$ (158)
@@
$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}\ \leq\ \alpha_{\ell,N}\ \leq\ \left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (160)
@@
$\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}$ $\ \leq\ \alpha_{\ell,N}-1\ \leq\ \sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{j}.$ (161)
@@
$\ =\ \frac{2N^{2\delta}}{\sqrt{k(N)}}\cdot\frac{1-\left(\frac{2N^{2\delta}}{\sqrt{k(N)}}\right)^{N^{\delta}}}{1-\frac{2N^{2\delta}}{\sqrt{k(N)}}},$ (162)
@@
$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/5}}{N^{\epsilon/2}}\ =\ 2N^{-3\epsilon/10}.$ (163)
@@
$\ =\ O\left(N^{-3\epsilon/10}\right).$ (164)
@@
$\left|\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (165)
Applying (A) and (165) to (A), we get
$\left|\alpha_{\ell,N}-1\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (166)
Thus, substituting the estimate (166) back to (A) gives us
$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-3\epsilon/10}\right).$ (167)
@@ Appendix B Case for $k_{S}<d$
$D_{1}(y)$ $\ :=\ \sum_{S\subsetneq[d]}\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)$
$\quad\quad\cdot\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}dz_{m-1}\cdots dz_{m+1-d}.$ (168)
@@
$\quad\quad\quad\cdot\left(\prod_{i\in[k_{S}+1;d-1]}\varphi(z_{m+i-d})\right)\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right).$ (169)
@@
$\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
@@
$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
$\quad\quad\quad\cdot A(z_{m+k_{S}-d})\cdot\left(\prod_{i=k_{S}+1}^{d-1}\varphi(z_{m+i-d})\right)\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (170)
To bound (B), we first give an estimate on the following
@@
$\quad\ \ll\ \frac{1}{\sqrt{2\pi(d-k_{S})}}e^{-\frac{\left(y-\sum_{i=1}^{k_{S}}z_{m+i-d}\right)^{2}}{2(d-k_{S})}},$ (171)
@@
$\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{k_{S}-1}\int_{z_{m}+(j-1)-d}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
@@
$\quad\ \ll\ N^{-1/2-\delta}\underbrace{\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\cdots\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}}_{k_{S}-1}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
@@
$\quad\ \ll\ N^{-1/2-\delta}\prod_{i\in S\setminus[k_{S}+1;d]}\left(\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\varphi(z_{m+i-d})dz_{m+i-d}\right)$
$\quad\quad\quad\cdot\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}\left(\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}A(z_{m+i-d})dz_{m+i-d}\right)$
$\quad\ \ll\ N^{-1/2-\delta+\left|[d]\setminus(S\cup\{k_{S}\})\right|(-1/2-\delta+\delta^{\prime})}\ \ll\ N^{-1/2-\delta},$ (172)
@@
Since there is only a finite number of proper subsets $S$ of $[d]$, then when $k_{S}<d$, we have that $D_{1}(y)\ll N^{-1/2-\delta}$ by definition of $D_{1}(y)$ in (B).
@@ Appendix C Case for $d>2$
In this section, we want to obtain the following estimate on $dI_{N}(y)/dy$ when $d>2$
$\frac{d}{dy}I_{N}(y)$
@@
$\quad\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (174)
@@
$\frac{d}{dy}I_{N}(y)$
$\quad\ \ll\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
$\quad\quad\quad+\Bigg(\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg)dz_{m-2}\cdots dz_{m+1-d}.$ (175)
Since $y$ is an upper bound on $\sum_{i=1}^{d}z_{m+i-d}$ and $z_{m-2}\leq z_{m-1}\leq z_{m}$, then we have $z_{m-2}\leq z_{m-1}\leq(z_{m-1}+z_{m})/2\leq(y-\sum_{i=1}^{d-2}z_{m+i-d})/2$. When $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi((y-\sum_{i=1}^{d-2}z_{m+i-d})/2)$. Hence,
@@
$\quad\ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1},$ (176)
where in the last line we use the change of variable $z_{m-1}\mapsto-z_{m-1}+y-\sum_{i=1}^{d-2}z_{m+i-d}$. The integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard Gaussian random variable, which is finite. Hence, when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$,
@@
$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right).$ (177)
@@
$y-\sum_{i=1}^{d-1}z_{m+i-d}\ \geq\ y-\sum_{i=1}^{d-2}z_{m+i-d}-\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\ =\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\ \geq\ 0.$ (178)
@@
$\quad\ =\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)-\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (179)
@@
$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)+\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (180)
@@
$\frac{d}{dy}I_{N}(y)$
$\quad\ \ll\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
@@
$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d},$ (181)